# Video In-context Learning

### v1

@@ Appendix A Additional Results
Appendix B Implementation Details
B.1 Model architectures
We utilize LLaMA as the Transformer architecture in our work. To validate the scaling behavior, we train three different sizes of the Vid-ICL model within the LLaMA architecture: 300M, 700M, and 1.1B. We carefully tune the hidden dimension, MLP intermediate dimension, and the number of Transformer decoder layers to achieve different model sizes. The configuration details of these models are presented in Table 8.
@@ B.2 Hyperparameters
The hyperparameters used to train the Vid-ICL model are presented in Table 9. We utilize inverse square root scheduler and start model training with 10000 warmup steps.
@@ B.3 Dataset statistics
The summary of dataset used in our paper is presented in Table 10. We provide detailed statistics including number of videos, total of tokens after tokenization and tokens used to train our Vid-ICL model.
Appendix C Broader Impact
Positive societal impacts of Vid-ICL include its potential to enhance various downstream applications such as embodied planning and robotic simulation by enabling models to imitate actions demonstrated in videos. As for negative societal impacts, the widespread adoption of this technique may raise concerns about privacy when powerful models gain the ability to analyze and generate video content at scale.
@@ Appendix D Ethics Statement
Vid-ICL is exclusively a research initiative with no current plans for product integration or public access. We are committed to adhering to Microsoft AI principles during the ongoing development of our models. The datasets utilized in this study are publicly available and have been thoroughly reviewed to ensure they do not include personally identifiable information or offensive content. Nonetheless, as these datasets are sourced from the Internet, there may still be inherent biases. To address this, we have implemented a rigorous filtering process on the training data to minimize the potential for the model to generate inappropriate content.

### v2

@@ Appendix B Implementation Details
B.1 Model Architecture
We utilize LLaMA as the Transformer architecture in our work. To validate the scaling behavior, we train three different sizes of the VidIT model within the LLaMA architecture: 300M, 700M, and 1.1B. We carefully tune the hidden dimension, MLP intermediate dimension, and the number of Transformer decoder layers to achieve different model sizes. The configuration details of these models are presented in Table 7.
@@ B.2 Hyperparameters
The hyperparameters used to train the VidIT model are presented in Table 8. We utilize inverse square root scheduler and start model training with 10,000 warmup steps.
@@ B.3 Dataset Statistics
The summary of dataset used in our paper is presented in Table 9. We provide detailed statistics including number of videos, total of tokens after tokenization and tokens used to train our VidIT model.
B.4 Training Details
Our largest variant, VidIT 1.1B, is trained on 2x8H100 nodes, with Pytorch DDP parallel strategy integrated in the Pytorch-Lightening trainer. The training time is about 3 hours per epoch and thus the total training GPU hours is 720 (15 epochs).
Appendix C Limitations
Despite the promising results, several limitations exist in this study. Firstly, while VidIT excels at generating short video sequences, its performance on longer sequences has not been thoroughly evaluated. Secondly, the model’s dependence on the quality and relevance of provided demonstrations can lead to inconsistencies, particularly when demonstrations are noisy or semantically misaligned. Future work should address these limitations by utilizing visual tokenizers with temporal compression, and scaling or developing robust models to handle imperfect demonstrations.
@@ Appendix D Ethics Statement
Positive societal impacts of VidIT include its potential to enhance various downstream applications such as embodied planning and robotic simulation by enabling models to imitate actions demonstrated in videos. As for negative societal impacts, the widespread adoption of this technique may raise concerns about privacy when powerful models gain the ability to analyze and generate video content at scale. The datasets utilized in this study are publicly available and have been thoroughly reviewed to ensure they do not include personally identifiable information or offensive content. Nonetheless, as these datasets are sourced from the Internet, there may still be inherent biases. To address this, we have implemented a rigorous filtering process on the training data to minimize the potential for the model to generate inappropriate content.
Table 10: More samples of VidIT generated samples on Something-something v2 dataset.
Demonstration Query & Generation
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Table 11: More examples of VidIT generated samples on RT-1 dataset.
Demonstration Query & Generation
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]